A quartz crystal vibrates with a frequency of 51,621 Hz. What is the period of the crystal's motion?
Answer #1

Given: Frequency (f) = 51621 Hz

Since. Period (T) = 1 / f = 1 / (51621 Hz) = 1.9372 x 10-5 sec

or Period (T) = 19.37 \mus